OpenStudy (anonymous):

solve the differential equation using the boundary condition: dy/dx = (x^2+x+1)/x, y=0 when x=1

OpenStudy (anonymous):

\[\frac{dy}{dx} = \frac{x^2 + x + 1}{x} \implies dy = (x + 1 + \frac{1}{x})dx\]Then you can integrate both sides.

OpenStudy (anonymous):

Then plug in 1 for x and solve for c1 :-)

OpenStudy (anonymous):

ohhhhhh i see thank you sooo much, i was making more difficult than it is lol thanks

OpenStudy (anonymous):

No problem :-)